10 November 2009 Lab Meeting
- Presentation: Dr Shichu Huang from Auburn. 2pm, PLEASE BE ON TIME.
- LUNCH at 12pm with Dr. Huang. Cathie's Treat. Please meet in at the 7th floor elevators. We will go to Bertucci's.
- Oakridge deadline is 1 Feb 2010.
- PAPERS Drafted before MicroTAS Are we done with this? 1st draft by Thur
‡ Flu R01:Integration
- New design - address evaporation loss, Reagent delivery and storage solution. Consider marginating flows?
- New design sent to make Rubber mold
† Sample Concentration (Lead: Jane, Team: Jaephil)
- MDCK Cells are not propagating. Try troubleshooting efforts. Sonali will do on Monday.
- Cassidy needs to see plaque assay again - so when cells are up, she needs training.
- Up to high 10^7 for positive control. Silver substrate no signal.
- Jane working on the cell lysate control.
- Main loss is at the outlet/sample collection: tangential filtration design in drawing stage.
- Contacted EC Shaw about rubber stamp mold for new design for evaporation.
- Set up COMSOL and StarCD, look for governing equations for simulation of evaporation.
- Committee meeting setup
† SPE Column Optimization for RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)
- Video with two color dyes and [red(primer) then blue, green, water+tween]
- repeat experiment reported on 10/27 with 700nm Silica 3X.
- controls for these: empties, non-silica monolith and full silica monolith, std. recipe.
- Load DNA in a straw of 5,50,100ul SPE, quick test of hypothesis: more SPE = more DNA, To be repeated with Alex and Mark
- Look into modifying Alex's pneumatic setup for the PATH guys. Instead of building the straw array over here NOTE: Sean
- PLanning net silica per channel experiments. with 700nm, 3x possible. Deliver around 11/5/09.
- New virus prep (JC, 6/11/09), getting Ct's of 27 for 1:16 dilution and 28 for 1:64 dilution. Use more diluted sample for future optimization experiments? Channels were 700nm silica/1X. pfu/ml is 10^6.
- Tried 2 more chips with 3X 0.15um Silica. Chips are breaking b/c of no crosslinking. SPE was coming out even at 2ml/hr flow rate.
- Running no silica and no SPE chips with virus. (JC/RNA) - No Silica channels grab comparable amounts of RNA with the channels with silica, lower with no SPE.
- PCR is running on the samples from the new channel design. Will compare results with Hussam after the meeting and report back if significantly different.
† PCR - CMI (Lead: Qingqing)
- QQ will work on the initial integration steps of SPE + RT (reservoir)+ PCR.
- third design was tested with water,big bubbles observed, but the water still could be collected at the outlet.
- third design was tested with PCR reagent(Flu assay). On-chip PCR does not work.
- the same experiment was repeated with the reduced hot-start time(five minutes). the on-chip PCR still does not work.
* New channel has been designed.
- molds(SU-8,PDMS,Epox)has been made.
- chip has been made and tested with water. Both thermal and fluidic control are fine. it will be tested with PCR reagent this week.
- primer for toxin A.
- on chip PCR with the optimized primer and template concentration does not work.
- MgCl2 concentration was optimized to improve the PCR efficiency.
- on chip PCR with the optimized MgCl2 concentration does not work.
- Primer for toxin B.
New primer sequence from Lisa has been ordered, will be tested when they come.
- PCR2 Paper formatted for LOAC this week.
need to get back from Cathie
- CMK: PCR 1 draft. Analytical Chem. MCK is running more simulations.
† HDA (Lead: Jaephil, Team:Sonali)
- Start planning R01 for Submission on 2/5/10. MM, JD, CMK.
- HDA chip fabrication training will be after new process settled (comeback from mTAS).
- Two methods, Hot embossing and cutter plotting will be combined to reduce whole process time. New mold design sent.
- Paper submitted 10/6.
‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )
- Sonali to train Lisa on SPE.
- QQ to run test PCR on chip with genomic DNA and Toxin B primers.
‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)
- Meeting 9am Monday (every other week)
‡ Agilent Automated Sample Preparation (Lead: Alex)
- Bacillus subtilis gDNA and cells were tested on SNAP2 machine. gDNA gave good results, pcr signals from cell-samples still low. More tests ongoing. Hot Dog samples will follow.
- First Draft of Yeast Paper send to Cathie and Alexis.
- Nano-C: which cells shall we use for lysing experiments?
‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)
- Metal piece modification for the integration - Fix did not turned out good. Make new or try fix again?
- paper 1, 1st draft by Thu(10.29): Evap with Sol Gel substrate. Not integrated. MSSA, E coli.
‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
- New experiments happening now.
- Cathie will submit paper inquiry.
‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)
- IRB approved.
- Cathie submitted the companion BU IRB form for exemption. Revised, still still waiting.
‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME), Suma)
- Phone meeting with PATH this week. 10/13.
- Senior Project Proposals due in less than two weeks! (23 Nov.)
- Frank: 1-week unstabilized sample PCRed (increased degradation with higher storage temps as expected; eluting 2-week sample on Wed.
- Frank: PVA to be ordered for home-made polymer storage solution; begin solubility tests upon arrival.
- Sean: Drawings with measurements have been made. Currently working on learning CAD software and making a professional CAD drawing.
- Sean met with Alex for an integration of straws on black machine. They sketched a design. Sean will make a professional drawing and create a required parts list. Alex will supervise the design. Before handing in the job to the shop floor, the design will be presented to Cathie.
- Mark to begin initial DNA testing with shorter time periods (1,3 days, etc.) to be completed next week. PCR training to be completed with Hussam.